# Charge transport and mode transition in dual-energy electron beam diodes

### v1

@@ Abstract
This Letter uncovers five distinct charge transport modes and their transitions in dual-energy electron beam diodes. We, via fully kinetic particle-in-cell (PIC) simulations, establish that the specific mode (e.g., space charge oscillations) and the current transport characteristics are essentially governed by the interplay between the electron beam energy and injected current density. A more generalized analysis is conducted for nn–component electron beams, and a theoretical piecewise function is proposed, which agrees well with the PIC results under designed conditions. The discovery provides a mechanistic picture of multiple electron beam transport in diodes, paving the way for novel designs of high-performance modern vacuum electronic devices.
@@
Introduction—Electron beams Tsimring (2007); Go (2022); Sharma and Hara (2025), as directed electron streams characterized by their current density Li et al. (2025); Campanell et al. (2012); Campanell and Umansky (2016); Jensen (2019), energy spectrum Furman and Pivi (2002); Chen and Fu (2025a) and spatiotemporal evolutions Heri and Zhang (2025); Ang and Zhang (2007), are indispensable drivers for various applications Zhang et al. (2017); Pagán Muñoz et al. (2024); Dubrovski et al. (2024); Chen and Fu (2025b), such as virtual cathode oscillators Jiang (2024); Mumtaz et al. (2024), thermionic energy converters Sitek et al. (2021); Abdul Khalid et al. (2016); Lin et al. (2024), compact terahertz sources Pedersen et al. (2010); Zhang et al. (2025); Kumar et al. (2025), and ultrafast switches Samizadeh Nikoo et al. (2020). In electron beam driven diodes, an injected intense electron current, accompanied by significant space charge effects Birdsall and Bridges (1966); Campanell et al. (2025), can limit the transmitted current to a maximum, which is known as the space-charge-limited (SCL) current $J_{\text{SCL}}$ Child (1911); Jaffé (1944). The charge transport in diodes with mono-energetic electron beams has been well understood, and the maximum transmitted current obeys the prediction by the SCL current $J_{\text{SCL}}$ Zhu et al. (2013, 2023); Garner et al. (2025). Previously, first-principle particle-in-cell (PIC) simulations determined that in mono-energetic electron beam driven diodes, the transmitted current $J_{\text{tran}}$ is as a function of the injected current $J_{\text{inj}}$, i.e., $J_{\text{tran}}=J_{\text{inj}}$ for $J_{\text{inj}}\leq J_{\text{SCL}}$ whereas $J_{\text{tran}}$ becomes saturated (or oscillating with nonzero emission velocity) for $J_{\text{inj}}>J_{\text{SCL}}$ Lafleur (2020, 2022). If $J_{\text{inj}}>J_{\text{SCL}}$, a virtual cathode (adjacent to the cathode) can form with an oscillating potential barrier, which periodically reflects less energetic electrons released from the cathode, thereby inducing space-charge oscillations Bridges and Birdsall (1963); Lin and Fu (2025); Zhu and Kong (2023). In the oscillation regime, existing analytical solutions can predict the spatial profiles of diode physical quantities (e.g., electric potential) Lafleur (2020). The scaling law for the space charge oscillation frequency has been theoretically derived and validated via PIC simulations Lin et al. (2025). However, the SCL current requires modification when electrons exhibit distributed energies (e.g., the Maxwell distribution for electron thermal motion) or multi-energy electron beams can form, and then the space charge oscillations can be damped, resulting in a new steady state Lafleur (2020). Although the SCL current and oscillatory dynamics in mono-energy electron beam diodes have been well studied, electron beams with discrete energy components (e.g., dual-energy electron beams) exhibit emergent complexities, remaining to be explored for an uncharted theoretical framework.
@@
In this Letter, we report five complete charge transport modes and their transitions in diodes driven by precisely controlled dual-energy electron beams via PIC simulations. We demonstrate the control of five distinct operation modes by independently modulating high- and low-energy electron beams, and identify that the velocity ratio dictates mode transitions, whereas the injected current ratio determines transmitted current scaling characteristics.
@@
Physical model—The schematic of the planar vacuum diode with cathode-anode gap distance of $d_{\text{gap}}$ driven by dual-energy electron beams is shown in Fig. 1. Two groups of electrons are injected from the cathode, including low-energy electrons $\text{e}_{1}$ at an initial velocity $v_{1}=\beta_{1}\sqrt{2eV/m_{e}}$ with an injected current density $J_{1}$ and high-energy electrons $\text{e}_{2}$ at an initial velocity $v_{2}=\beta_{2}\sqrt{2eV/m_{e}}$ with an injected current density $J_{2}$, where $\beta_{1}$ and $\beta_{2}$ are dimensionless initial velocity factors, $e$ is the elementary charge, $V$ is the gap voltage, and $m_{e}$ is the electron mass. Here, the initial velocity and injected current of each electron beam can be independently controlled, which are characterized by two dimensionless parameters, i.e., the velocity ratio $k_{1}=v_{1}/v_{2}=\beta_{1}/\beta_{2}\in(0,1)$ and the injected current ratio $k_{2}=J_{1}/J_{0}\in(0,1)$, where $J_{0}=J_{1}+J_{2}$ is the total injected current density.
@@
The virtual cathode with the minimum potential $\phi_{\text{vc}}$ at the position $x_{\text{vc}}$ can be generated self-consistently by both electron beams, which can block electrons when their initial energy $\mathcal{E}_{0}<-e\phi_{\text{vc}}$. Note that here we have $\mathcal{E}_{0}=\mathcal{E}_{e1}$ for $\text{e}_{1}$, $\mathcal{E}_{0}=\mathcal{E}_{e2}$ for $\text{e}_{2}$, and $\mathcal{E}_{e1}<\mathcal{E}_{e2}$. Thus, electrons with different energies exhibit distinct dynamical behaviors responding to the electric potential. As shown in Fig. 1, due to the time-dependent electric potential, low-energy electrons $\text{e}_{1}$ without sufficient kinetic energy are predominantly reflected; whereas high-energy electrons $\text{e}_{2}$ can overcome potential barriers, which are most transported and absorbed by the anode. The electron reflection point can be determined by the critical condition where the reflection potential $\phi_{\text{ref}}=-\mathcal{E}_{0}/e$, and here we have $\phi_{\text{ref,1}}=-\mathcal{E}_{e1}/e$ for $\text{e}_{1}$ and $\phi_{\text{ref,2}}=-\mathcal{E}_{e2}/e$ for $\text{e}_{2}$. Thus, the electron reflection point position for low-energy electrons is closer to the cathode than for high-energy electrons, i.e., $x_{\text{ref,1}}<x_{\text{ref,2}}$.
@@
Figure 1: Schematic of the dual-energy electron beam driven diode, in which the cathode injection current consists of low-energy and high-energy electrons. Low-energy electrons tend to be reflected mostly, whereas high-energy electrons can be transmitted predominantly to the anode. Note that the cathode-to-virtual cathode region is enlarged and $x_{\text{ref,1}}<x_{\text{ref,2}}<x_{\text{vc}}\ll d_{\text{gap}}$, where $x_{\text{ref,1}}$ and $x_{\text{ref,2}}$ are the reflection points, $x_{\text{vc}}$ is the virtual cathode position, and $d_{\text{gap}}$ is the gap distance.
@@
Operation modes—Each electron beam (low-energy $\text{e}_{1}$ and high-energy $\text{e}_{2}$) exhibits three possible charge transport states: fully transported mode (T–mode), oscillation mode (O–mode) with intermittent transport and reflection, and fully reflected mode (R–mode). While state permutations permit $3^{2}=9$ theoretical transport modes, stronger space-charge suppression on $\text{e}_{1}$ than $\text{e}_{2}$ constrains the system to five dynamical modes (M1–M5). Notably, both $\text{e}_{1}$ and $\text{e}_{2}$ cannot be fully reflected simultaneously; otherwise, the virtual cathode is unable to sustain. By systemically tuning the control parameters of each electron beam, e.g., initial velocity and injected current density, we identify five observationally distinct operation modes in dual-energy electron beam driven diodes, as shown in Table 1.
@@
Figure 2 demonstrates five operation modes (M1–M5), including the electric potential distribution $\phi-x$, temporal evolution of the cathode surface electric field $E_{\text{s}}-t$, and phase space of spatially averaged electron density versus cathode surface electric field $\overline{n}_{e}-E_{\text{s}}$ for each mode. In mode M1, the minimum gap potential follows $\phi_{\text{min}}>\phi_{\text{ref,1}}>\phi_{\text{ref,2}}$ [see Fig. 2(a1)], which enables fully transport of both low-energy electrons $\text{e}_{1}$ and high-energy electrons $\text{e}_{2}$ in steady state. Mode M2 emerges when the intense space-charge effect occurs and lowers the electric potential, forming a virtual cathode potential $\phi_{\text{vc}}$ that oscillates near the reflection potential of low-energy electrons $\phi_{\text{ref,1}}$. This condition, though maintain full transport for $\text{e}_{2}$, can trigger intermittent reflection-transport cycles for $\text{e}_{1}$, which can induce oscillations of physical quantities, such as the cathode surface electric field $E_{\text{s}}$ [Fig. 2(b2)]. Since the oscillation is induced by $\text{e}_{1}$, the oscillation frequency for M2 can be predicted by $f_{\text{os}}\varpropto{J}_{1}^{1/2}{\beta}_{1}^{-1/2}$ Lin et al. (2025). Mode M3 is established when the virtual cathode potential stabilizes below $\phi_{\text{ref,1}}$ but above $\phi_{\text{ref,2}}$, i.e., $\phi_{\text{ref,1}}>\phi_{\text{min}}>\phi_{\text{ref,2}}$, [Fig. 2(a3)], resulting in a new steady state where $\text{e}_{1}$ is completely reflected but $\text{e}_{2}$ is fully transported. In Fig. 2(a4), by further increasing the injected current, M4 occurs when the virtual cathode potential is sufficiently low to oscillate near the reflection potential of high-energy electrons $\phi_{\text{ref,2}}$. Thus, $\text{e}_{1}$ can be completely reflected and reabsorbed by the cathode whereas $\text{e}_{2}$ has intermittent transport and reflection, inducing periodic oscillation with the oscillation frequency scaling as $f_{\text{os}}\varpropto J_{2}^{1/2}{\beta}_{2}^{-1/2}$. Especially, when the energies of $\text{e}_{1}$ and $\text{e}_{2}$ are rather close (e.g., $k_{1}=\sqrt{\mathcal{E}_{e1}/\mathcal{E}_{e2}}=0.95$), M5 is formed and the virtual cathode potential oscillates between $\phi_{\text{ref,1}}$ and $\phi_{\text{ref,2}}$ [Fig. 2(a5)]. Then both the low- and high-energy electrons can oscillate at their respective frequencies, generating a combined oscillation waveform [see Fig. 2(b5)]. The phase space distributions of $\overline{n}_{e}-E_{\text{s}}$ for five operation modes (M1–M5) are shown in Figs. 2(c1)-(c5), respectively, which confirm that loop trajectories are formed in the oscillation modes.
@@
Figure 3: The total transmitted current $J_{\text{tran}}$ ($J_{\text{tran}}=J_{\text{tran,1}}+J_{\text{tran,2}}$ with $J_{\text{tran,1}}$ for low-energy electron $\text{e}_{1}$ and $J_{\text{tran,2}}$ for high-energy electron $\text{e}_{2}$) and reflected current $J_{\text{ref}}$ ($J_{\text{ref}}=J_{\text{ref,1}}+J_{\text{ref,2}}$ with $J_{\text{ref,1}}$ for $\text{e}_{1}$ and $J_{\text{tran,2}}$ for $\text{e}_{2}$) as a function of the injected current $J_{0}$ ($J_{0}=J_{1}+J_{2}$ with $J_{1}$ for $\text{e}_{1}$ and $J_{2}$ for $\text{e}_{2}$) under different conditions of two electron beams. (a) Initial velocity ratio $k_{1}=\beta_{1}/\beta_{2}=0.1$ and injected current percentage $k_{2}=J_{1}/J_{0}=0.5$ for $\text{e}_{1}$; (b) $k_{1}=0.5$ and $k_{2}=0.8$; (c) $k_{1}=0.5$ and $k_{2}=0.2$; (d) $k_{1}=0.95$ and $k_{2}=0.5$. (e) Schematic of the operation mode transition for panel (a). (f)–(g) Illustrations of the transmitted current for panels (b) and (c). (h) Schematic of the operation mode transition for panel (d).
Transmitted current characteristics—The transmitted current characteristics for mono-energetic electrons have been established and rigorously verified in previous studies Lafleur (2020). The SCL current, as the critical current threshold Lafleur (2020), is expressed as
@@
where $\beta$ considers the initial velocity of emitted electrons. The transmitted current $J_{\text{tran}}$ equals the injected current $J_{\text{inj}}$ for $J_{\text{inj}}\leq J_{\text{SCL}}$ while becomes saturatee for $J_{\text{inj}}>J_{\text{SCL}}$. However, in dual-energy electron beam driven diodes, each electron beam possesses its own SCL current ($J_{\text{SCL,1}}$ for $\text{e}_{1}$ and $J_{\text{SCL,2}}$ for $\text{e}_{2}$) since $\beta$ is equal to $\beta_{1}$ or $\beta_{2}$, respectively, which leads to significantly complex transmitted current behaviors.
@@
By separately tuning the control parameters for dual-energy electron beams, the transmitted current scaling and charge transport mode transition are determined, as shown in Fig. 3. The ratio of initial velocity for two electron beams, $k_{1}=\beta_{1}/\beta_{2}$, plays a significant role in determining the charge transport mode transition. For electron beams with a large difference in velocity (energy) (e.g., $k_{1}=0.1$), the transmitted current characteristics are illustrated in Fig. 3(a). When the injected current $J_{0}$ is sufficiently low (below the SCL current of the low-energy electrons $J_{\text{SCL,1}}$), both low- and high-energy electrons can be fully transmitted [mode M1]. When the injected current for low-energy electrons $J_{1}$ exceeds their corresponding SCL current $J_{\text{SCL,1}}$, the transport of low-energy electrons $\text{e}_{1}$ is suppressed while the high-energy electrons $\text{e}_{2}$ can still be transported completely [mode M2]. Notably, in this regime, the transmitted current for low-energy electrons $J_{\text{tran,1}}$ falls below $J_{1}$ and the reflected current for low-energy electrons $J_{\text{ref,1}}$ increases substantially, maintaining the current conservation $J_{1}=J_{\text{tran,1}}+J_{\text{ref,1}}$. As the injected current further increases, low-energy electrons can be fully reflected with $J_{\text{ref,1}}=J_{\text{inj,1}}$, while high-energy electrons remain fully transported [mode M3]. Once the injected current for high-energy electrons $J_{2}$ is larger than their corresponding SCL current $J_{\text{SCL,2}}$, high-energy electrons $\text{e}_{2}$ are partially reflected, accompanied with the entire reflection of $\text{e}_{1}$ [mode M4]. Consequently, while increasing the total injected current $J_{0}$, the charge transport modes undergo a well-defined pathway: $\text{M}1\rightarrow\text{M}2\rightarrow\text{M}3\rightarrow\text{M}4$, which can be abstracted to discrete eigen-states defined by combinatorial permutations of the transport modes for both energy components [as illustrated in Fig. 3(e)].
@@
Figure 3(b) maps the transmitted current characteristics under parameter conditions of $k_{1}=\beta_{1}/\beta_{2}=0.5$ and $k_{2}=J_{1}/J_{0}=0.8$, demonstrating more universal behaviors of the transmitted currents. Note that the critical mode transitions are quantitatively defined by the intersection points where the injected current of each electron beam component reaches its corresponding SCL current. As illustrated in the schematic diagram [see Fig. 3(f)], we identify several critical points (A-D) to visually mark the evolution of transmitted current. At point A, the total injected current $J_{0}=J_{\text{SCL,1}}$ and all the electrons can be transported steadily, i.e., M1 with the total transmitted current $J_{\text{tran}}=J_{0}$. The transition from M1 to M2 occurs abruptly when the injected current of low-energy electrons reaches their SCL current [point B], i.e., $J_{1}=J_{\text{SCL,1}}$, and then $J_{\text{tran}}$ deviates from $J_{0}$. Theoretically, the transition point is expected when $J_{1}<J_{\text{SCL,1}}$; however, due to the significant energy difference between the two electron beams and the relatively minor space charge effect of the high-energy electrons, one can consider the critical transition at point B. The point C marks the M2–M3 transition where $J_{\text{tran}}=J_{2}$ and $J_{\text{tran,1}}=0$, indicating that only the single electron beam ($\text{e}_{2}$) is transmitted in the diode. At point D, $J_{2}=J_{\text{SCL,2}}$ and the transmitted current achieves its maximum, simultaneously signifying the M3–M4 transition.
@@
In addition, the injected current ratio $k_{2}=J_{1}/J_{0}$ serves as a key control parameter, modulating the mode transition pathway. Figure 3(c) shows the transmitted current characteristics under low-energy electron-dominated injection conditions ($k_{2}=0.8>0.5$), which shows an obvious difference from that in high-energy electron-dominated injection conditions ($k_{2}=0.2<0.5$) [see Fig. 3(b)]. Notably, low-energy electrons $\text{e}_{1}$ undergo abnormal reflection before reaching their conventional SCL current $J_{\text{SCL,1}}$, leading to an effective SCL current $J_{\text{SCL,1}}^{\text{eff}}$ smaller than $J_{\text{SCL,1}}$. This is because the virtual cathode induced by space charge effects, dominated by high-energy electron injection, easily suppresses the transport of low-energy electrons. As illustrated in Fig. 3(g), the critical point B for M1–M2 transition is defined as the intersection point with $J_{1}=J_{\text{SCL,1}}^{\text{eff}}$, which occurs before reaching the conventional threshold $J_{1}=J_{\text{SCL,1}}$.

### v2

@@ Abstract
This Letter uncovers five distinct charge transport modes and their transitions in dual-energy electron beam diodes. We via first-principle particle-in-cell (PIC) simulations establish that the specific mode (e.g., space charge oscillations) and the current transport characteristics are essentially governed by the interplay between the electron beam energy and injected current density. A generalized analysis is conducted for nn–component electron beams, and a theoretical piecewise function is for the transmitted current density proposed, which agrees well with the PIC results under designed conditions. The discovery provides a mechanistic picture of multiple electron beam transport in diodes, paving the way for novel designs of high-performance modern vacuum electronic devices.
@@
Introduction—Electron beams Tsimring (2007); Go (2022); Sharma and Hara (2025), as directed electron streams characterized by their current density Li et al. (2025); Campanell et al. (2012); Campanell and Umansky (2016), energy spectrum Wenz et al. (2019); Furman and Pivi (2002); Chen and Fu (2025a) and spatiotemporal evolutions Heri and Zhang (2025); Ang and Zhang (2007), are indispensable drivers for various applications Zhang et al. (2017); Pagán Muñoz et al. (2024); Dubrovski et al. (2024); Chen and Fu (2025b), such as virtual cathode oscillators Jiang (2024); Mumtaz et al. (2024), thermionic energy converters Sitek et al. (2021); Abdul Khalid et al. (2016); Lin et al. (2024), compact terahertz sources Pedersen et al. (2010); Zhang et al. (2025); Kumar et al. (2025), and ultrafast switches Samizadeh Nikoo et al. (2020). In electron beam driven diodes, an injected intense electron current, accompanied by significant space charge effects Birdsall and Bridges (1966); Campanell et al. (2025), can limit the transmitted current to a maximum, which is known as the space-charge-limited (SCL) current $J_{\text{SCL}}$ Child (1911); Jaffé (1944). The charge transport in diodes with mono-energetic electron beams has been well understood, and the maximum transmitted current obeys the prediction by the SCL current $J_{\text{SCL}}$ Zhu et al. (2013, 2023); Garner et al. (2025). Previously, first-principle particle-in-cell (PIC) simulations showed that in mono-energetic electron beam driven diodes, the transmitted current $J_{\text{tran}}$ is as a function of the injected current $J_{\text{inj}}$, i.e., $J_{\text{tran}}=J_{\text{inj}}$ for $J_{\text{inj}}\leq J_{\text{SCL}}$ whereas $J_{\text{tran}}$ becomes saturated (or oscillating with nonzero emission velocity) for $J_{\text{inj}}>J_{\text{SCL}}$ Lafleur (2020, 2022). If $J_{\text{inj}}>J_{\text{SCL}}$, a virtual cathode (adjacent to the cathode) can form with an oscillating potential barrier, which periodically reflects less energetic electrons released from the cathode, thereby inducing space-charge oscillations Bridges and Birdsall (1963); Lin and Fu (2025); Zhu and Kong (2023). In the oscillation regime, existing analytical solutions can predict the spatial profiles of diode physical quantities (e.g., electric potential) Lafleur (2020). The scaling law for the space charge oscillation frequency has been theoretically derived and validated via PIC simulations Lin et al. (2025). However, the SCL current fails when electrons exhibit distributed energies (e.g., the Maxwell distribution for electron thermal motion) or multi-energy electron beams can form, since the space charge oscillations can be damped, resulting in a new steady state Lafleur (2020). Although the SCL current and oscillatory dynamics in mono-energy electron beam diodes have been well studied, electron beams with discrete energy components (e.g., dual-energy electron beams) exhibit emergent complexities, remaining to be understood for an uncharted theoretical framework.
@@
In this Letter, we report five complete charge transport modes and their transitions in diodes driven by precisely controlled dual-energy electron beams. For the first time, we, via PIC simulations, demonstrate the control of five distinct operation modes by independently modulating high- and low-energy electron beams, and identify that the velocity ratio dictates mode transitions, whereas the injected current ratio determines transmitted current scaling characteristics.
@@
Physical model—The schematic of the planar vacuum diode with cathode-anode gap distance of $d_{\text{gap}}$ driven by dual-energy electron beams is shown in Fig. 1. Two groups of electrons are injected from the cathode, including low-energy electrons $\text{e}_{1}$ at an initial velocity $v_{1}=\beta_{1}\sqrt{2eV/m_{e}}$ with an injected current density $J_{1}$ and high-energy electrons $\text{e}_{2}$ at an initial velocity $v_{2}=\beta_{2}\sqrt{2eV/m_{e}}$ with an injected current density $J_{2}$, where $\beta_{1}$ and $\beta_{2}$ are dimensionless initial velocity factors, $e$ is the elementary charge, $V$ is the gap voltage, and $m_{e}$ is the electron mass. The different groups of electron beams, generated from field emission Forbes (2019); Ying et al. (2025), thermal-field emission Jensen (2019), photo-emission Heimerl et al. (2024), or laser-driven accelerators Wenz et al. (2019), can be independently transported. The total injected current density $J_{0}$ is the sum of injected current density of $\text{e}_{1}$ and $\text{e}_{2}$, i.e., $J_{0}=J_{1}+J_{2}$. Here, the initial velocity and injected current of each electron beam can be independently controlled, which are characterized by two dimensionless parameters, i.e., the velocity ratio $v_{1}/v_{2}=\beta_{1}/\beta_{2}\in(0,1)$ and the injected current ratio $J_{1}/J_{2}$ (can be controlled greater or less than one).
@@
The virtual cathode with the minimum potential $\phi_{\text{vc}}$ at the position $x_{\text{vc}}$ is generated self-consistently by both electron beams, which can block electrons when their initial energy $\mathcal{E}_{0}<-e\phi_{\text{vc}}$. Here we have $\mathcal{E}_{0}=\mathcal{E}_{e1}$ for $\text{e}_{1}$, $\mathcal{E}_{0}=\mathcal{E}_{e2}$ for $\text{e}_{2}$, and $\mathcal{E}_{e1}<\mathcal{E}_{e2}$, thus, electrons with different energies exhibit distinct dynamical behaviors responding to the electric potential. As shown in Fig. 1, low-energy electrons $\text{e}_{1}$ without sufficient kinetic energy can be predominantly reflected by the virtual cathode potential barrier, whereas high-energy electrons $\text{e}_{2}$ can overcome potential barriers, and can be mostly transported and absorbed by the anode. The electron reflection point is determined by the critical condition where the reflection potential $\phi_{\text{ref}}=-\mathcal{E}_{0}/e$, and here we have $\phi_{\text{ref,1}}=-\mathcal{E}_{e1}/e$ for $\text{e}_{1}$ and $\phi_{\text{ref,2}}=-\mathcal{E}_{e2}/e$ for $\text{e}_{2}$. Thus, the electron reflection point position for low-energy electrons is closer to the cathode than for high-energy electrons, i.e., $x_{\text{ref,1}}<x_{\text{ref,2}}$.
@@
Figure 1: Schematic of the dual-energy electron beam driven diode, in which the cathode injection current consists of low-energy and high-energy electrons. Low-energy electrons tend to be reflected mostly, whereas high-energy electrons can be transmitted predominantly to the anode. $x_{\text{c}}=0$ is the cathode, and $x_{\text{a}}=d_{\text{gap}}$ is the anode. Note that the cathode-to-virtual cathode region is enlarged and $x_{\text{ref,1}}<x_{\text{ref,2}}<x_{\text{vc}}\ll d_{\text{gap}}$, where $x_{\text{ref,1}}$ and $x_{\text{ref,2}}$ are the reflection points, and $x_{\text{vc}}$ is the virtual cathode position.
@@
Operation modes—Each electron beam (low-energy $\text{e}_{1}$ and high-energy $\text{e}_{2}$) exhibits three possible charge transport states: fully transported mode (T–mode), oscillation mode (O–mode) with intermittent transport and reflection, and fully reflected mode (R–mode). While state permutations permit $3^{2}=9$ theoretical transport modes, stronger space-charge suppression on $\text{e}_{1}$ than $\text{e}_{2}$ constrains the system to five dynamical modes (M1–M5). Notably, both $\text{e}_{1}$ and $\text{e}_{2}$ cannot be fully reflected simultaneously; otherwise, the virtual cathode is unable to sustain Lafleur (2020). By systemically tuning the control parameters of each electron beam, e.g., initial velocity and injected current density, we identify five observationally distinct operation modes in dual-energy electron beam driven diodes, as shown in Table 1. The "NA" indicates the combination does not exist.
@@
Figure 2 demonstrates five operation modes (M1–M5), including the electric potential distribution $\phi-x$, temporal evolution of the cathode surface electric field $E_{\text{s}}-t$, and phase space of spatially averaged electron density versus cathode surface electric field $\overline{n}_{e}-E_{\text{s}}$ for each mode. In mode M1, the minimum gap potential follows $\phi_{\text{min}}>\phi_{\text{ref,1}}>\phi_{\text{ref,2}}$ [see Fig. 2(a1)], which enables fully transport of both low-energy electrons $\text{e}_{1}$ and high-energy electrons $\text{e}_{2}$ in steady state. Mode M2 emerges when the intense space-charge effect occurs and lowers the electric potential, forming a virtual cathode potential $\phi_{\text{vc}}$ that oscillates near the reflection potential of low-energy electrons $\phi_{\text{ref,1}}$. This condition, though maintain full transport for $\text{e}_{2}$, can trigger intermittent reflection-transport cycles for $\text{e}_{1}$, which can induce oscillations of physical quantities, such as the cathode surface electric field $E_{\text{s}}$ [Fig. 2(b2)]. Since the oscillation is induced by $\text{e}_{1}$, the oscillation frequency for M2 can be predicted by $f_{\text{os}}\varpropto\sqrt{J_{1}/\beta_{1}}$ Lin et al. (2025). Mode M3 is established when the virtual cathode potential stabilizes below $\phi_{\text{ref,1}}$ but above $\phi_{\text{ref,2}}$, i.e., $\phi_{\text{ref,1}}>\phi_{\text{min}}>\phi_{\text{ref,2}}$, [Fig. 2(a3)], resulting in a new steady state where $\text{e}_{1}$ is completely reflected but $\text{e}_{2}$ is fully transported. In Fig. 2(a4), by further increasing the injected current, M4 occurs when the virtual cathode potential is sufficiently low to oscillate near the reflection potential of high-energy electrons $\phi_{\text{ref,2}}$. Thus, $\text{e}_{1}$ can be completely reflected and reabsorbed by the cathode whereas $\text{e}_{2}$ has intermittent transport and reflection, inducing periodic oscillation with the oscillation frequency scaling as $f_{\text{os}}\varpropto\sqrt{J_{2}/\beta_{2}}$. Especially, when the energies of $\text{e}_{1}$ and $\text{e}_{2}$ are rather close (e.g., $v_{1}/v_{2}=\sqrt{\mathcal{E}_{e1}/\mathcal{E}_{e2}}=0.95$), M5 is formed and the virtual cathode potential oscillates between $\phi_{\text{ref,1}}$ and $\phi_{\text{ref,2}}$ [Fig. 2(a5)]. Then both the low- and high-energy electrons can oscillate at their respective frequencies, generating a combined oscillation waveform [see Fig. 2(b5)]. The phase space distributions of $\overline{n}_{e}-E_{\text{s}}$ for five operation modes (M1–M5) are shown in Figs. 2(c1)-(c5), respectively, which confirm that loop trajectories are formed in the oscillation modes.
@@
Figure 3: Total transmitted current $J_{\text{tran}}$ ($J_{\text{tran}}=J_{\text{tran,1}}+J_{\text{tran,2}}$ with $J_{\text{tran,1}}$ for low-energy electron $\text{e}_{1}$ and $J_{\text{tran,2}}$ for high-energy electron $\text{e}_{2}$) and reflected current $J_{\text{ref}}$ ($J_{\text{ref}}=J_{\text{ref,1}}+J_{\text{ref,2}}$ with $J_{\text{ref,1}}$ for $\text{e}_{1}$ and $J_{\text{tran,2}}$ for $\text{e}_{2}$) as a function of the injected current $J_{0}$ ($J_{0}=J_{1}+J_{2}$ with $J_{1}$ for $\text{e}_{1}$ and $J_{2}$ for $\text{e}_{2}$) under different conditions. Governing parameters ($\beta_{1}/\beta_{2},\ J_{1}/J_{2}$)= (0.1, 1) for (a), (0.5, 4) for (b), (0.5, 0.25) for (c), and (0.95, 1) for (d). (e) Schematic of the operation mode transition for panel (a). (f)–(g) Illustrations of the transmitted current for panels (b) and (c). (h) Schematic of the operation mode transition for panel (d).
Transmitted current characteristics—The transmitted current characteristics for mono-energetic electrons have been explicitly demonstrated in previous studies Lafleur (2020); Lin et al. (2025). The SCL current, as the critical current threshold Lafleur (2020), is expressed as
@@
where $\beta$ quantifies the initial velocity of emitted electrons. The transmitted current $J_{\text{tran}}$ equals the injected current $J_{\text{inj}}$ for $J_{\text{inj}}\leq J_{\text{SCL}}$ while becomes saturatee for $J_{\text{inj}}>J_{\text{SCL}}$. However, in dual-energy electron beam driven diodes, each electron beam possesses its own SCL current ($J_{\text{SCL,1}}$ for $\text{e}_{1}$ and $J_{\text{SCL,2}}$ for $\text{e}_{2}$) since $\beta$ is equal to $\beta_{1}$ or $\beta_{2}$, respectively, which leads to significantly complex transmitted current behaviors.
@@
By separately tuning the control parameters for dual-energy electron beams, the transmitted current scaling and charge transport mode transition are determined, as shown in Fig. 3. The ratio of initial velocity for two electron beams, $\beta_{1}/\beta_{2}$, plays a significant role in determining the charge transport mode transition. For electron beams with a large difference in velocity (energy) (e.g., $\beta_{1}/\beta_{2}=0.1$), the transmitted current characteristics are illustrated in Fig. 3(a). When the injected current $J_{0}$ is sufficiently low (below the SCL current of the low-energy electrons $J_{\text{SCL,1}}$), both low- and high-energy electrons can be fully transmitted [mode M1]. When the injected current for low-energy electrons $J_{1}$ exceeds their corresponding SCL current $J_{\text{SCL,1}}$, the transport of low-energy electrons $\text{e}_{1}$ is suppressed while the high-energy electrons $\text{e}_{2}$ can still be transported completely [mode M2]. Notably, in this regime, the transmitted current for low-energy electrons $J_{\text{tran,1}}$ falls below $J_{1}$ and the reflected current for low-energy electrons $J_{\text{ref,1}}$ increases substantially, maintaining the current conservation $J_{1}=J_{\text{tran,1}}+J_{\text{ref,1}}$. As the injected current further increases, low-energy electrons can be fully reflected with $J_{\text{ref,1}}=J_{\text{inj,1}}$, while high-energy electrons remain fully transported [mode M3]. Once the injected current for high-energy electrons $J_{2}$ is larger than their corresponding SCL current $J_{\text{SCL,2}}$, high-energy electrons $\text{e}_{2}$ are partially reflected, accompanied with the entire reflection of $\text{e}_{1}$ [mode M4]. Consequently, while increasing the total injected current $J_{0}$, the charge transport modes undergo a well-defined pathway: $\text{M}1\rightarrow\text{M}2\rightarrow\text{M}3\rightarrow\text{M}4$, which can be abstracted to discrete eigen-states defined by combinatorial permutations of the transport modes for both energy components [as illustrated in Fig. 3(e)].
@@
Figure 3(b) maps the transmitted current characteristics under parameter conditions of $\beta_{1}/\beta_{2}=0.5$ and $J_{1}/J_{2}=4$, demonstrating more universal behaviors of the transmitted currents. Note that the critical mode transitions are quantitatively defined by the intersection points where the injected current of each electron beam component reaches its corresponding SCL current. As illustrated in the schematic diagram [see Fig. 3(f)], we identify several critical points (A-D) to visually mark the evolution of transmitted current. At point A, the total injected current $J_{0}=J_{\text{SCL,1}}$ and all the electrons can be transported steadily, i.e., M1 with the total transmitted current $J_{\text{tran}}=J_{0}$. The transition from M1 to M2 occurs abruptly when the injected current of low-energy electrons reaches their SCL current [point B], i.e., $J_{1}=J_{\text{SCL,1}}$, and then $J_{\text{tran}}$ deviates from $J_{0}$. Theoretically, the transition point is expected when $J_{1}<J_{\text{SCL,1}}$; however, due to the significant energy difference between the two electron beams and the relatively minor space charge effect of the high-energy electrons, one can consider the critical transition at point B. The point C marks the M2–M3 transition where $J_{\text{tran}}=J_{2}$ and $J_{\text{tran,1}}=0$, indicating that only the single electron beam ($\text{e}_{2}$) is transmitted in the diode. At point D, $J_{2}=J_{\text{SCL,2}}$ and the transmitted current achieves its maximum, simultaneously signifying the M3–M4 transition.
@@
In addition, the injected current ratio $J_{1}/J_{2}$ serves as a key control parameter, modulating the mode transition pathway. Figure 3(c) shows the transmitted current characteristics under low-energy electron-dominated injection conditions ($J_{1}/J_{2}=4>1$), which shows an obvious difference from that in high-energy electron-dominated injection conditions ($J_{1}/J_{2}=0.25<1$) [see Fig. 3(b)]. Notably, low-energy electrons $\text{e}_{1}$ undergo abnormal reflection before reaching their conventional SCL current $J_{\text{SCL,1}}$, leading to an effective SCL current $J_{\text{SCL,1}}^{\text{eff}}$ smaller than $J_{\text{SCL,1}}$. This is because the virtual cathode induced by space charge effects, dominated by high-energy electron injection, easily suppresses the transport of low-energy electrons. As illustrated in Fig. 3(g), the critical point B for M1–M2 transition is defined as the intersection point with $J_{1}=J_{\text{SCL,1}}^{\text{eff}}$, which occurs before reaching the conventional threshold $J_{1}=J_{\text{SCL,1}}$.